### v3

# Diversity of Low-mass Planet Atmospheres in the C–H–O–N–S–Cl System with Interior Dissolution, Nonideality, and Condensation: Application to TRAPPIST-1e and Sub-Neptunes

Abstract
A quantitative understanding of the nature and composition of low-mass rocky (exo)planet atmospheres during their evolution is needed to interpret observations. The magma ocean stage of terrestrial- and sub-Neptune planets permits mass exchange between their interiors and atmospheres, during which the mass and speciation of the atmosphere is dictated by the planet’s volatile budget, chemical equilibria, and gas/fluid solubility in molten rock. As the atmosphere cools, it is modified by gas-phase reactions and condensation. We combine these processes into an open-source Python package built using JAX called Atmodeller, and perform calculations for planet sizes and conditions analogous to TRAPPIST-1e and K2-18b. For TRAPPIST-1e-like planets, our simulations indicate that CO-dominated atmospheres are prevalent during the magma ocean stage, which, upon isochemical cooling, predominantly evolve into CO2-rich atmospheres of a few hundred bar at 280 K. Around 40% of our simulations predict the coexistence of liquid water, graphite, -sulfur, and ammonium chloride—key ingredients for surface habitability. For sub-Neptune gas dwarfs, pressures are sufficiently high (GPa) that gas fugacities deviate from ideality, thereby drastically enhancing solubilities. This buffers the total atmospheric pressure to lower values than for the ideal case. These effects conspire to produce CH4-rich sub-Neptune atmospheres for total pressures exceeding 3.5 GPa, provided H/C is 100 solar and O2 moderately reducing (3 log10 units below the iron-wüstite buffer). Otherwise, molecular hydrogen remains the predominant species at lower total pressures and/or higher H/C. For all planets at high temperature, solubility enriches C/H in the atmosphere relative to the initial composition.
1 Introduction
Whilst the discovery of exoplanets has proliferated in recent years, a significant challenge lies in unraveling how their bulk interior compositions relate to atmospheric speciation. Based on current demographics studies, the most abundant kinds of planets are those with masses in between those of Earth and Neptune and are divided into ”super-Earths,” which are rocky planets with radii less than 1.5 , and ”sub-Neptunes,”’ which are planets with radii larger than 2 that likely have H2-rich atmospheres (Bean et al., 2021). Characterizing planets with possible molten surfaces (magma oceans)—such as close-in rocky and super-Earth planets in addition to sub-Neptunes with thick hydrogen (H2) envelopes—demands a comprehensive treatment of thermodynamics at the melt-atmosphere interface (Chao et al., 2021; Charnoz et al., 2023; Misener et al., 2023; Rigby et al., 2024; Seo et al., 2024; Werlen et al., 2025) because the mass of the melt reservoirs always predominates over that contained in the atmosphere, such that the physicochemical state of the liquid (i.e., the chemical potentials of components set by pressure, temperature, and bulk composition) controls the near-surface atmospheric speciation, thereby setting a lower boundary condition for the atmosphere (Hirschmann, 2012; Sossi et al., 2020). Above this boundary, the atmospheric speciation may be modified by dynamics, chemistry, photochemistry, and atmospheric escape (Catling & Kasting, 2017; Wordsworth & Kreidberg, 2022; Selsis et al., 2023; Krissansen-Totton et al., 2024; Owen & Schlichting, 2024; Rogers et al., 2024; Cherubim et al., 2025; Nicholls et al., 2025). Ultimately, the spectral signal in the uppermost atmosphere is most accessible for interrogation by space-based observatories such as the James Webb Space Telescope (JWST). Hence, adequately modellng physical and chemical processes at the melt-atmosphere interface is foundational to interpreting current and upcoming observations.
During planet formation, volatiles delivered in planetary building blocks (e.g., Alexander et al., 2012; Marty, 2012) and by capture of the nebular gas (e.g., Sharp & Olson, 2022; Young et al., 2023) likely equilibrated with the growing planet’s rocky magma ocean, whose capacity to store volatile species varies according to their solubilities. A solubility law for any given volatile species relates its dissolved concentration in magma to its fugacity, which is, in turn, related to partial pressure through a fugacity coefficient. These laws are determined either experimentally or theoretically via ab initio simulations for a set of conditions, including total pressure, temperature, oxygen fugacity, and melt composition, that affect the stable melt and/or gas species. Solubility studies have focused on deriving relations for major atmosphere-forming species (e.g., H2O, H2, CO2, CO, CH4, N2, SO2, Cl2, and the noble gases) in silicate melts (e.g., Dixon et al., 1995; Newcombe et al., 2017; Sossi et al., 2023; Hirschmann et al., 2012; Ardia et al., 2013; Libourel et al., 2003; Boulliung & Wood, 2022; Bernadou et al., 2021; O’Neill & Mavrogenes, 2022; Thomas & Wood, 2021; Davis & Caracas, 2024; Jambon et al., 1986), and were largely conducted at moderate to high-pressure conditions (hundreds of bar to several GPa) since they were originally devised to cover the conditions of Earth’s mantle and crust. These conditions, however, are also comparable to the melt-atmosphere interface of large rocky planets and hence these solubility relations are now finding new applications in exoplanet research.
For interface pressures greater than 1 GPa, gases can no longer be reasonably treated as obeying the ideal gas law, depending on temperature and the nature of the gas species (Bridgman, 1924; Duan & Zhang, 2006). Fugacity coefficients quantify how strongly a gas departs from ideal behavior and are often expressed in terms of temperature (), pressure (), and composition () (e.g., Redlich & Kwong, 1949; Tian & Heng, 2024). Deviations from unity pertain physically to differences in volume at a given pressure and temperature relative to the ideal gas law, which are products of (1) the finite size of gas particles and (2) attractive/repulsive interactions between them (Van der Waals, 1873) and are themselves a function of , and . They can be derived from real gas equations of state (EOS) (e.g., Saxena & Fei, 1988; Holland & Powell, 1991), thermodynamic tables and charts (e.g., Chase, 1998), or directly from experimental data and measurements (e.g., Presnall, 1969; Jakobsson & Oskarsson, 1994; Frost & Wood, 1997). Nonideality is of direct relevance to interpreting the formation and evolution of rocky planets, which are mostly rock by mass and gas by volume. Both gas-phase reactions and gas solubilities are regulated by fugacities, which are only equivalent to partial pressures in the ideal limit. Meanwhile, pressure (not fugacity) is the operative quantity in satisfying volatile mass balance (e.g., Appendix C, Bower et al., 2022). In essence, nonideality demands that ”thermodynamic pressure,” relevant to reactions and solubility, and ”mechanical pressure,” relevant to mass balance, are treated as distinct entities. This may lead to exoplanet demographics that are notably different from those derived by assuming ideal gas behavior and chemically inert rocky cores (e.g., Kite et al., 2019).
For planets that subsequently cool and crystallize their mantles, such as the solar system bodies, volatiles exsolve to form an early outgassed atmosphere that sets the initial state for the subsequent long-term evolution of the atmosphere (e.g., Bower et al., 2019, 2022; Nikolaou et al., 2019; Katyal et al., 2020; Lichtenberg et al., 2021; Gaillard et al., 2021; Liggins et al., 2022; Krissansen-Totton et al., 2024; Maurice et al., 2024). The gaseous species in these atmospheres, in addition to any condensed phases that form upon cooling, may be the earliest ingredients necessary for prebiotic chemical reactions to take place (Sossi et al., 2020; Zahnle et al., 2020; Benner et al., 2020; Preiner & Martin, 2021; Johansen et al., 2024; Korenaga, 2025). However, no systematic study of how volcanically degassed atmospheres cool and condense under a range of conditions distinct from that of the Earth has been undertaken. Therefore, whether the raw chemical ingredients for the origin of life were endemic to Earth, and to what extent they occur in other planetary systems, remains unanswered. Moreover, super-Earths and sub-Neptunes may sustain a molten interior, either due to close proximity to their host star or an insulating envelope, and hence volatiles could remain sequestered in the molten interior, giving rise to a wide variety of atmospheres. Hence, quantifying both the volatile budget of rocky worlds that potentially become habitable (e.g., Earth and Earth twins), as well as the interior–atmosphere exchange occurring on super-Earths and sub-Neptunes, requires compositionally consistent and accurate volatile solubility relations applicable to the unique conditions at the interior–atmosphere interface.
We have therefore devised a versatile and modular toolkit (Atmodeller) that the community can utilize and extend to compute volatile partitioning between planetary reservoirs with solubility, nonideality, and condensation. In particular, this will enable new theoretically and experimentally derived solubility laws and fugacity coefficients to be efficiently and accurately incorporated into models devised by (exo)planetary scientists. Ultimately, Atmodeller provides a transparent and robust framework for planetary outgassing calculations. After presenting the theory and method on which Atmodeller is based (Section 2), we then showcase its features by calculating the earliest atmosphere of TRAPPIST-1e (Section 3.1) as well as the atmosphere of a gas dwarf sub-Neptune (Section 3.2).
2 Method
2.1 Formulation and numerical solution
We focus on providing an overview of the specific design considerations of Atmodeller, noting that a review of the general approaches to formulate and solve chemical equilibrium calculations are presented in Leal et al. (2017). Gas and (optionally) condensed species are specified by the user and are used to determine an independent reaction set via Gaussian elimination (see Appendix A, Leal et al., 2016). Atmodeller then applies the extended law of mass-action (xLMA) equations (Leal et al., 2016) to create a closed set of nonlinear equations that combines equilibrium chemistry with elemental mass balance, cast as a root-finding problem. The xLMA has an intuitive formulation using equilibrium constants and allows for the simultaneous solution of all (user-specified) species in the system without a priori knowledge of the species’ thermodynamic stability, which is particularly important to facilitate solutions with condensates (see also Kitzmann et al., 2024). Solubility laws and real gas EOS provide optional closure conditions, where, if desired, solubility can be neglected and the system assumed ideal. In this regard, Atmodeller can also model planetary atmospheres above chemically inert mantles or cool solidified surfaces; there is no inherent requirement for magma to be present nor the gas and condensed matter to chemically interact.
If included, solubility relations ensure chemical equilibrium between a species’ fugacity and its dissolved abundance in melt (Section 2.2), whereas real gas EOS relate a species’ partial pressure to its fugacity (Section 2.3). The mass balance is formulated by considering elemental abundances in condensates, in the molten mantle, and in the atmosphere (e.g., Appendix A, Bower et al., 2019), which requires the specification of some planetary properties (e.g., surface temperature, mantle mass, melt fraction, etc.). The mantle mass (and hence the gravitational acceleration on the planet) remains constant, irrespective of the mass of volatiles dissolved in it. However, insofar as the masses of the combined volatile elements added never exceed 3%–4% of the total planetary mass, even in the case of the most volatile-rich sub-Neptunes (Section 3.2), this approximation introduces minimal error. Hence, in this formulation, the mass of volatiles in the atmosphere directly determines the magma-atmosphere surface pressure, and therefore is inextricably tied to the partitioning of gaseous species between the atmosphere and molten mantle. Atmodeller allows the specification of mixed constraints, such that fugacity constraints can be combined with elemental mass constraints as long as the system of equations is closed. This facilitates a common, desirable scenario in which oxygen fugacity is defined, in absolute terms as O2, or relative to a mineralogical buffer (e.g., the iron-wüstite buffer; IW) while other elements have fixed abundance.
Atmodeller is written in the Python programming language for seamless integration with existing modeling frameworks and leverages JAX (Bradbury et al., 2018) for high-performance computation. It consists of three subpackages that contain JAX-compliant solubility laws, real gas EOS, and thermodynamic data, all of which can be accessed independently and used as components of Python-based modeling frameworks if desired (with or without JAX support). Notably, the automatic differentiation feature of JAX is convenient for its ability to propagate derivative information through solubility laws and real gas EOS, which often present complex functional forms that may themselves require iterative solution during the overall solution procedure. Another feature, automatic vectorization, enables Atmodeller to easily spawn batch calculations to efficiently explore parameter space in a single vectorized solve. Finally, just-in-time compilation enables a model to be pre-compiled, which facilitates fast Atmodeller calculations as part of a repeating workflow such as a time integration in evolution models. Atmodeller uses Optimistix (Rader et al., 2024) to solve the root-finding problem and additionally leverages Lineax for linear solves (Rader et al., 2023) and Equinox (Kidger & Garcia, 2021). As with all systems of nonlinear equations, obtaining a solution requires that applied constraints (fugacity, mass) can be satisfied simultaneously with the given system of equations (mass balance, equilibrium chemistry) and closure conditions (solubility laws, free energies of individual gas species, real gas EOS). We independently verified the results of Atmodeller by comparison to FactSage 8.2 (Bale et al., 2016) calculations assuming ideal gas behavior (Appendix C), and the Python package has additional comparisons with condensates as well as FastChem 3.1.1 (Kitzmann et al., 2024). Furthermore, each solubility law and real gas EOS implemented in Atmodeller includes an associated test that verifies its output against tabulated data or published figures from the original source, ensuring the accuracy and reproducibility of the model components.
2.2 Solubility laws
| Experimental Calibration† | |||||
| Species | Composition∗ | Reference | Pressure | Temperature | O2 rel. IW |
| (kbar) | (K) | (log10 units) | |||
| H2 | Basalt, Andesite (Fe-free) | Hirschmann et al. (2012, their Table 2) | 7–30 | 1673–1773 | -1 to 3.8 |
| H2 | Silicate glass (obsidian) | Gaillard et al. (2003, their Table 4) | 0.001–0.265 | 1073 | -0.36 to 0.6 |
| H2O | Basalt | Dixon et al. (1995, their Fig. 4) | 0.2–0.72 | 1473 | 4.2 to 5.5 |
| H2O | Peridotite | Sossi et al. (2023) | 0.001 | 2173 | -1.9 to 6.0 |
| H2O | Lunar basalt, An-Di | Newcombe et al. (2017, their Fig. 5) | 0.001 | 1623 | -3.0 to 4.8 |
| H2O | Basalt | Mitchell et al. (2017, their Fig. 8) | 10 | 1473 | 4.2 |
| He | Basalt (tholeiitic) | Jambon et al. (1986) | 0.001 | 1523–1873 | 7.3 to 10.7 |
| Ne | Basalt (tholeiitic) | Jambon et al. (1986) | 0.001 | 1523–1873 | 7.3 to 10.7 |
| Ar | Basalt (tholeiitic) | Jambon et al. (1986) | 0.001 | 1523–1873 | 7.3 to 10.7 |
| Kr | Basalt (tholeiitic) | Jambon et al. (1986) | 0.001 | 1523–1873 | 7.3 to 10.7 |
| Xe | Basalt (tholeiitic) | Jambon et al. (1986) | 0.001 | 1523–1873 | 7.3 to 10.7 |
| CO | Basalt, Rhyolite | Yoshioka et al. (2019) | 2–30 | 1473–1773 | 1 to 3.8 |
| CO | Basalt | Armstrong et al. (2015, their Eq. 10) | 10–12 | 1673 | -3.65 to 1.46 |
| CO2 | Basalt | Dixon et al. (1995, their Eq. 6) | 0.21–0.98 | 1473 | 4.2 to 5.5 |
| CH4 | Basalt (Fe-free) | Ardia et al. (2013, their Eq. 7a and 8) | 7–30 | 1673–1723 | -9.50 to -1.36 |
| N2 | Basalt (tholeiitic) | Libourel et al. (2003, their Eq. 23) | 0.001 | 1673–1698 | -8.3 to 8.7 |
| N2 | Basalt | Dasgupta et al. (2022) | 0.001–82 | 1323–2600 | -8.3 to 8.7 |
| N2 | Basalt | Bernadou et al. (2021) | 0.8–10 | 1473–1573 | -4.7 to 4.9 |
| S2‡ | Basalt, Andesite | Boulliung & Wood (2022, 2023) | 0.001 | 1473–1773 | -0.14 to 10.9 |
| Cl2 | Basalt, An-Di-Fo | Thomas & Wood (2021, their Fig. 4) | 15 | 1673 | 2.03 |
Note. — ∗Anorthite (An), Diopside (Di), Forsterite (Fo). : Range of experimental conditions used to determine the solubility laws. The oxygen fugacity, O2 is given in log10 units relative to the IW buffer. : Combines the solubility of sulfur as sulfide (S2-) and sulfate (SO/S6+).
The atmospheric compositions of planets with molten interiors are modulated by both the volatile endowment of the planet and the solubility of these volatiles in the melt. The solubility of a gas species depends on its fugacity (or partial pressure under ideal gas conditions) in the gas phase, the composition of the melt, and sometimes additional parameters, such as the oxygen fugacity (O2), temperature, and total pressure. Solubility laws for major atmosphere-forming species are determined from experiments that, so far, have largely been focused on Earth-like melt compositions and conducted under , conditions motivated by Earth’s mantle. These empirically determined laws often take the form of power-law or linear fits to the experimental data to relate the dissolved concentration of a species in the melt to its fugacity in the gas phase. Importantly, solubility experiments are performed for certain melt compositions under specific temperature and pressure conditions, meaning that extrapolating these laws beyond their calibrated conditions can be erroneous. The composition of the gas phase also affects the solubility by expanding or shrinking the stability of both gas and melt species in the system. For example, the dissolution of N may occur as NH, NH3, CN-, N3- and N2 moieties in silicate melts depending on the fugacities of N-, C- and H-bearing species at a given , (Dalou et al., 2019; Grewal et al., 2020), such that solubilities are not independent of atmospheric composition. Atmodeller includes a library of solubility laws from the literature for H2, H2O, CO, CO2, CH4, N2, S2, Cl2 and the noble gases He, Ne, Ar, Kr, Xe for various melt compositions and chemical systems (Table 1). We note the need for more solubility experiments of major volatiles for diverse melt compositions beyond those of the modern Earth, and under a wide range of , conditions as well as gas-phase compositions.
2.3 Real gas equations of state
Atmodeller includes a library of real gas EOS (Table 2), which yield expressions for the volume of any given species as a function of pressure and temperature. With this information, it is possible to quantify the relationship between a species’ fugacity and its partial pressure, expressed as the fugacity coefficient , where is fugacity and is pressure:
where is the gas constant, temperature, molar volume, , and the compressibility factor . The lower bound of the integration is the standard state pressure (usually 1 bar or 1 atmosphere, depending on the thermodynamic data source), and conventionally defines an ideal state, i.e. . Typically, experimental data or numerical calculations determine or , which are then fit to an empirical or physically motivated function of temperature and pressure, such as the compensated Redlich-Kwong EOS (CORK, Holland & Powell, 1991) or the virial EOS (Shi & Saxena, 1992). Because the behavior of gases under high is related to their critical temperature and pressure, often, the pressure and temperature dependence of volume is expressed relative to these critical quantities (, ), i.e., the law of corresponding states (Van der Waals, 1873). This enables regression of experimental data across similarly behaving species and thereby improves the average fit. For known pressure and temperature, fugacity coefficients and fugacities can then be computed by evaluating the integral in Equation 1. Some functional forms have convenient closed-form solutions of the fugacity, whereas others require numerical solution (see Holland & Powell, 1991, for a discussion). It is also beneficial for numerical stability to ensure the EOS are bounded and smooth, so we extrapolate the compressibility factor, , linearly according to a virial-type expression (Appendix D).
| Range† | |||||
| Species | EOS∗ | Reference | Note | Pressure | Temperature |
| (kbar) | (K) | ||||
| Ar | Virial | Saxena & Fei (1987a) | Corresponding states | 0-1000 | 200-3000 |
| CH4 | CORK | Holland & Powell (1991) | Corresponding states | 0-50 | 400-1900 |
| CH4 | Virial | Shi & Saxena (1992) | Corresponding states | 0-1000 | 200-3000 |
| CH4 | B-B | Holley et al. (1958) | 0-1 | 200-500 | |
| CO | CORK | Holland & Powell (1991) | Corresponding states | 0-50 | 400-1900 |
| CO | Virial | Shi & Saxena (1992) | Corresponding states | 0-1000 | 200-3000 |
| CO2 | CORK | Holland & Powell (1991) | Corresponding states | 0-50 | 400-1900 |
| CO2 | CORK | Holland & Powell (1991) | 0-50 | 400-1900 | |
| CO2 | CORK | Holland & Powell (1998) | Updated virial coefficients | 0-120 | 400-1900 |
| CO2 | Virial | Shi & Saxena (1992) | Corresponding states | 0-1000 | 200-3000 |
| CO2 | B-B | Holley et al. (1958) | 0-0.1 | 300-1000 | |
| COS | Virial | Shi & Saxena (1992) | Corresponding states | 0-1000 | 200-3000 |
| H2 | CORK | Holland & Powell (1991) | Corresponding states | 0-50 | 400-1900 |
| H2 | Virial | Saxena & Fei (1987a) | Corresponding states | 0-1000 | 200-3000 |
| H2 | Virial | Shi & Saxena (1992) | Corresponding states | 0-1000 | 200-3000 |
| H2 | B-B | Holley et al. (1958) | 0-1 | 20-600 | |
| H2 | ab initio | Chabrier & Debras (2021) | Also mixtures with He | 0-1014 | 100- |
| H2O | CORK | Holland & Powell (1991) | 0-50 | 400-1700 | |
| H2O | CORK | Holland & Powell (1998) | Updated virial coefficients | 0-120 | 400-1700 |
| H2S | CORK | Holland & Powell (2011) | Corresponding states | 0-50 | 400-1900 |
| H2S | Virial | Shi & Saxena (1992) | 0-10 | 400-1100 | |
| He | B-B | Holley et al. (1958) | 0-1 | 200-400 | |
| He | ab initio | Chabrier & Debras (2021) | Also mixtures with H2 | 0-1014 | 100- |
| N2 | CORK | Holland & Powell (1991) | Corresponding states | 0-50 | 400-1900 |
| N2 | Virial | Saxena & Fei (1987a) | Corresponding states | 0-1000 | 200-3000 |
| N2 | B-B | Holley et al. (1958) | 0-0.1 | 100-1000 | |
| NH3 | B-B | Holley et al. (1958) | 0-0.1 | 300-500 | |
| O2 | Virial | Shi & Saxena (1992) | Corresponding states | 0-1000 | 200-3000 |
| O2 | B-B | Holley et al. (1958) | 0-0.1 | 100-1000 | |
| S2 | CORK | Holland & Powell (2011) | Corresponding states | 0-50 | 400-1900 |
| S2 | Virial | Shi & Saxena (1992) | Corresponding states | 0-1000 | 200-3000 |
| SO2 | Virial | Shi & Saxena (1992) | 0.5-10 | 500-1100 | |
Note. — ∗Compensated-Redlich Kwong (CORK), Beattie-Bridgeman (B-B). : Approximate range of thermodynamic conditions, accessed by experiments or ab initio calculations, used to inform the real gas EOS.
2.4 Condensation
The condition for condensation is met when the addition of a condensed phase, liquid or solid, to the system of gas species satisfies the xLMA equations. We recall that by construction the xLMA accommodates both stable and unstable condensates in the same framework through the specification of stability variables, which are independent for each condensed species and solved self-consistently as part of the solution process. For implementation details, the reader is referred to Leal et al. (2016) and Kitzmann et al. (2024). For numerical stability, it is necessary to define a parameter that controls the minimum amount that a stable species can achieve. Following Leal et al. (2016), we adopt , where is the minimum amount of an element in the system that composes a given species. In principle, stability variables could also be applied to gas species, although since gas-phase chemistry is notably sensitive to O2—even at very low abundance—introducing a minimum cut-off for the amounts of gas species could distort results.
In Atmodeller, only pure condensed phases are treated, and, as such, the phase is stable when its activity, . The activity is defined as the ratio of the fugacity of any component at , , relative to that of the pure phase at standard state conditions; and bar:
where, by definition, for a pure phase. We do not account for the compressibilities of condensed phases (i.e., the Poynting correction is negligible). The stable condensates included in the base package are H2O(l), C(cr), -S(cr), and ClH4N(cr,l)—ammonium chloride, because these are of primary importance for assessing planetary habitability. Thermodynamic data for condensates as well as gas species are sourced from McBride et al. (2002) and Chase (1998).
3 Results
3.1 TRAPPIST-1e and similar Earth-sized planets
3.1.1 Context and parameters
There is significant interest in probing the atmospheres of temperate Earth-sized rocky planets, such as the TRAPPIST-1 planets (Gillon et al., 2017), , to understand the nature of potentially habitable environments beyond Earth and the solar system. The earliest outgassed atmospheres form during and immediately following planet assembly, whereby one or several epochs of magma ocean formation is thought to have distributed volatile elements readily between the atmosphere, mantle, and, potentially the metallic core (e.g., Hirschmann, 2012; Lebrun et al., 2013; Sossi et al., 2020; Chen & Jacobson, 2022; Gu et al., 2024; Dasgupta et al., 2024; Huang et al., 2024). This establishes the initial budget of volatiles in the atmosphere and interior, which may subsequently participate in geological volatile cycling that further modulates their atmospheric abundances (e.g., Sleep & Zahnle, 2001; Krissansen-Totton et al., 2018; Hakim et al., 2021). Here, we apply Atmodeller to a young TRAPPIST-1e analog as an example of a small rocky planet during its magma ocean phase following iron-nickel-alloy core formation, although the results are applicable broadly for any Earth-sized rocky planet. These results are used to inform subsequent Atmodeller calculations to determine the postformation atmosphere and surface abundances of TRAPPIST-1e at its present-day equilibrium temperature (Section 3.1.3).
TRAPPIST-1e has a total mass of 4.13 kg, a mantle mass of 2.91 kg (assuming the same core-mantle mass proportion as Earth), and surface radius of 5861 km (0.9 R⊕) (Agol et al., 2021). To compute the properties of the high-temperature atmosphere, we assume a surface temperature of 1800 K, We consider 14 gas species: This encompasses the majority of the stable, gaseous molecules of the so-called CHONS group of elements, in addition to Cl, that constitute the majority of living matter. .
Of the gas species listed above, the following solubility laws are used to distribute them between the atmosphere and the magma ocean: H2O (Dixon et al., 1995), H2 (Hirschmann et al., 2012), CO (Yoshioka et al., 2019), CO2 (Dixon et al., 1995), CH4 (Ardia et al., 2013), N2 (Libourel et al., 2003), S-bearing species (Boulliung & Wood, 2022, 2023), Cl-bearing species (Thomas & Wood, 2021). Both the EOS of pure species and the mixing between them are assumed to be ideal. Condensed carbon (graphite) is also allowed to exist depending on thermodynamic stability, since this high-temperature condensate, if present, can buffer the fugacities of carbon-bearing species at the most reducing conditions.
We do not explicitly model the metallic core as a reservoir for volatiles, but rather calibrate the initial elemental abundances relative to the bulk silicate Earth (BSE), which is best constrained from natural samples (e.g., Palme & O’Neill, 2014). This allows us to model the exchange of volatiles between the silicate mantle and atmosphere without necessitating comparatively uncertain high , partitioning data that governs equilibrium between silicate and metallic phases. Nevertheless, our approach does not mandate that the metallic core is volatile free, but rather that it is isolated during the period of atmosphere-mantle exchange that we consider. Alternatively, the core could be thought of as volatile free (relatively, compared to the mantle), if the prescribed abundances are to be correlated 1:1 with the bulk composition.
Therefore, we fix the total abundances of N, S, and Cl based on BSE estimates, scaled to the mass of TRAPPIST-1e. The average BSE estimates for N, S, and Cl are 7.01 kg, 9.47 kg, and 1.09 kg, respectively (Sakuraba et al., 2021; Hirschmann, 2016; Kendrick et al., 2017), and, when scaled to the mantle mass of TRAPPIST-1e, they are 4.85 kg, 6.55 kg, and 7.57 kg, respectively. In a Monte Carlo simulation with 10,000 realizations we uniformly sample, in log10 space, the total amount of hydrogen between 0.1 and 10 Earth oceans , C/H (by mass) between 0.1 and 10, and the oxygen fugacity by 5 log10 units either side of the IW buffer (Appendix E). All simulation data are available to download (Bower, 2025).
3.1.2 Volatile partitioning during the magma ocean stage
We consider TRAPPIST-1e with a fully molten mantle, for which solubility and gas-phase equilibria are both important in determining the atmospheric mass and composition. For comparison, Appendix B presents the same analysis assuming a partially molten mantle with 10% melt fraction, with major and minor species shown in Figure B1 and Figure B2, respectively. For a fully molten mantle, solubility conspires to restrict the diversity of atmospheres (Figure 1) compared to the low melt fraction case (Figure B1). The data are colored according to the oxygen fugacity relative to the IW buffer (IW), which defines whether the atmosphere is reduced (lower values) or oxidized (higher values). The scatter plots capture the diversity of atmospheres, whereas the contoured density levels and marginal distributions reveal the most probable range of values . Figure 1 and Figure 2 show major and minor species, respectively.
Reduced atmospheres are dominated (volume mixing ratio, VMR ) by either CO at low total H budgets or by H2 at high total H budgets, because neither carbon monoxide nor molecular hydrogen are particularly soluble in silicate melts (Hirschmann et al., 2012; Yoshioka et al., 2019). Furthermore, H2 can be accompanied by modest amounts (up to 15 mol%) of CH4, where CH4 is buffered by graphite precipitation:
At O2s higher than the IW buffer, the mixing ratio of CH4 drops precipitously as it is mostly replaced by H2O and CO2, as described by the homogeneous equilibrium:
However, because H2O(g) is more soluble in silicate melt than H2(g) or CO2(g), it never becomes a dominant species in the atmosphere, even under oxidizing conditions (see also Bower et al., 2022; Sossi et al., 2023). Consequently, the abundances of hydrogen-bearing species in the atmosphere remain low, with H2O(g) and H2(g) comprising just a few percent of the total. Therefore, owing to its relatively lower solubility, CO2(g) assumes the role of the prevailing gas species in atmospheres between IW = +1 and IW = +3. At the highest O2s above IW = +3, SO2(g) overtakes CO2(g) as the prevailing gas species for BSE-like S abundances (see also Gaillard et al., 2022; Maurice et al., 2024; Gillmann et al., 2024). Because no other gaseous species form in appreciable quantities at these conditions, the mixing ratios of CO2(g) and SO2(g) are anticorrelated in oxidizing atmospheres.
The dominant outcome of preferential H2O dissolution is to enhance the relative abundances of carbon-bearing species in the atmosphere (Figure 3). This is illustrated by the mean molar mass () frequency distribution (Figure 1, bottom right), which shows a low plateau between the molar masses corresponding to H2 and H2O, a dominant peak around that of CO, a smaller peak around the molar mass of CO2, and a tail extending to SO2. The small peak at 7–8 g mol-1 corresponds to the most reducing, CH4–H2 atmospheres. For ”neutral” O2 atmospheres 1–2 log units either side of the IW buffer, the near-quantitative removal of H2O from the atmosphere means that peaks in the molar mass frequency distribution can be ascribed to a singular, dominant gas species (Figure 1, bottom row). Therefore, an atmosphere with g mol-1 must be CO-rich, rather than a mixture of lighter hydrogen-bearing and heavier carbon-bearing species, because H2O(g) is never a major gas species and CO2–H2 atmospheres are excluded thermodynamically (see also Sossi et al., 2023). The most oxidizing atmospheres have molar masses that are defined by binary mixtures of CO2 and SO2. As such, the mean molar mass of an atmosphere is broadly positively correlated with its O2, where atmospheres with g mol-1 must be both reducing and H-rich (Figure 4, bottom left).
A fully molten mantle always dissolves at least 60% of the total hydrogen budget. At IW close to , high C/H atmospheres have marginally less total H dissolved, all else being equal, due to the stability of CH4(g) over H2(g). For all but these most reduced conditions, the dissolved fraction of H approaches 100% (Figure 4). This is evident by the sharp peak around 100% in the marginal distribution of dissolved hydrogen where even the 10th percentile exceeds 97% (see also Sossi et al., 2020). As a result, atmospheric C/H ratios can be around 4 orders of magnitude larger than total C/H, with the disparity increasing to more oxidized conditions due to increasing H2O/H2. Nearly no C is dissolved in the mantle under reducing conditions (owing to the low solubilities of CO(g) and CH4(g)), but this value increases up to about 25% of the total carbon inventory at the most oxidized conditions where CO2(g) is stable (Figure 4). This proportion decreases for correspondingly lower melt fractions (Figure B3).
Total pressures range between 100 and 104 bar (90% lie within 5–500 bar), with atmospheres produced in equilibrium at intermediate O2s near the IW buffer delimiting both the high- and low ends of the range. The total pressure is largely governed by the number of ocean masses of H, together with the C/H ratio, with high C/H resulting in the highest atmospheric pressures at a given O2. Even though the mantles in equilibrium with reducing atmospheres store proportionally less H and C than their more oxidizing counterparts, their lower mean molecular masses () result in only modest total pressures (Equation F1). At high O2s near IW = +4, the increasing degree of CO2 dissolution into the mantle restricts the total pressure of SO2-rich atmospheres to around 200 bar. Sulfur is almost entirely dissolved below IW = +3, but above this value up to 60% of the sulfur inventory can exist in the atmosphere (Figure 4). The nonmonotonic solubility of S arises due to the occurrence of two moieties in silicate liquids; S2- below IW = +4 and SO above this value (O’Neill & Mavrogenes, 2022; Boulliung & Wood, 2023). Because the S content of the silicate liquid is proportional to O (for constant temperature, S2 and FeO content) when S2- is stable, and to O when SO is stable, the solubility of S exhibits a minimum at S2-/SO = 0.5 (where is mole fraction), or roughly IW = +4 (see also Hughes et al., 2023). Therefore, SO2 should decrease in atmospheres that are more oxidizing and more reducing than those formed at the solubility minimum.
Minor species for a 100% molten mantle are shown in Figure 2 (see Figure B2 for the 10% molten equivalent). Of these, N2(g) can become a major species (i.e., its mixing ratio exceeds 0.5) in a handful of cases. These atmospheres correspond to those with the lowest total pressures. That is, since the abundances of N, Cl, and S are held constant (while those of H, C, and O vary) in the Monte Carlo simulations, N2-rich atmospheres arise when the prescribed initial abundances of H and C are low. This behavior is also reflected in the inverse correlation between Cl2(g) mixing ratio (which never exceeds 1.510-7) and total pressure (Figure 2). Indeed, the mixing ratios of N2(g) and Cl2(g) are positively correlated. The other minor species diverge from this expectation because their partial pressures are sensitive to the presence of other volatile species, whose fugacities depend on O2 and/or the bulk chemical composition of the atmosphere.
The second most abundant minor species is S2(g), reaching mixing ratios up to 3% (0.03), with these cases corresponding exclusively to oxidizing atmospheres. As detailed above, more reducing conditions promote the dissolution of S as S2- in the silicate liquid, thereby restricting the S budget of the atmosphere and thus its ability to stabilize S2(g). A third S-bearing gas species, H2S(g), is less abundant than S2(g) and SO2(g) in atmospheres more oxidized than IW=+2, yet it becomes the dominant S-bearing gas species (mixing ratios up to 0.005) at O2s below this value. This reflects the increase in H2 at the expense of H2O, promoting the formation of H2S(g) by:
The higher mixing ratios of NH3(g) (up to 0.0001) and HCl(g) (up to 0.015) in reduced atmospheres relative to their oxidized counterparts (Figure 2) are explained by reactions analogous to Equation 5 occurring when substituting S2 for N2 and for Cl2, respectively. Since the abundance of HCl on a molar basis in our simulations, HCl(g) is the dominant Cl-bearing gas under all conditions. We note, however, the lack of experimentally determined solubilities for Cl, N, and S in H-bearing systems. The mixing ratios in the 10% molten cases show generally similar behavior, though the more soluble elements (Cl and S) exhibit higher atmospheric mixing ratios compared to those determined in the 100% molten cases, owing to the declining mass of molten material capable of storing them in the mantle. The corollary is that the more insoluble species (e.g., N2(g)) have relatively lower mixing ratios.
3.1.3 Cooled atmosphere and elemental surface inventory
The magma ocean stage establishes the initial volatile inventories in the atmosphere (Section 3.1.2), and as the atmosphere cools and condenses it distributes materials at the planetary surface that can participate in geochemical cycles and fuel a potential biosphere. To this end, we begin with the calculated atmospheric elemental abundances for each of the fully molten mantle cases presented in Section 3.1.2 and cool the atmospheres, isochemically (i.e., assuming a constant bulk atmosphere composition inherited from the magma ocean stage), to 280 K. This is a few kelvin above the freezing point of water at 1 bar and lies between the estimates of the equilibrium temperatures of TRAPPIST-1d and TRAPPIST-1e assuming a null albedo (Delrez et al., 2018). We include the contribution of any graphite present in the abundance of carbon of the high-temperature atmosphere, given that graphite remains close to the planetary surface (Keppler & Golabek, 2019) and in equilibrium with the atmosphere. In addition to the aforementioned gas species and graphite (Section 3.1.1), we further allow for the presence of liquid water (H2O), rhombic sulfur (also known as -sulfur), and ammonium chloride (ClH4N) depending on thermodynamic stability, since these condensed phases can be major carriers of elements in the system. The same analysis is applied to the 10% molten mantle cases in Appendix B.
| Dominant Species | Total Pressure (bar) | Count with Condensates | |||||
| (VMR 50%) | Count | Min | Max | Median | H2O | H2O+C | H2O+C+S+ClH4N |
| At 280 K, following equilibration with a fully molten mantle at 1800 K | |||||||
| H2 | 150 | 12.25 | 268.41 | 65.20 | 150 | 0 | 0 |
| CO2 | 8336 | 0.45 | 8991.13 | 58.73 | 7709 | 6549 | 4406 |
| CH4 | 546 | 0.21 | 436.59 | 32.96 | 546 | 491 | 0 |
| N2 | 92 | 0.11 | 2.38 | 1.89 | 40 | 40 | 1 |
| SO2 | 823 | 6.51 | 231.68 | 133.86 | 580 | 0 | 0 |
| ClH | 9 | 0.16 | 0.62 | 0.40 | 9 | 9 | 0 |
| None | 44 | 1.74 | 234.88 | 17.17 | 19 | 1 | 1 |
| Total | 10,000 | 9053 | 7090 | 4408 | |||
| At 280 K, following equilibration with a partially molten mantle at 1800 K | |||||||
| H2 | 1634 | 0.76 | 675.95 | 66.50 | 1634 | 0 | 0 |
| CO2 | 5230 | 0.53 | 9547.52 | 104.16 | 5230 | 4002 | 4002 |
| CH4 | 1403 | 0.23 | 1009.14 | 62.98 | 1403 | 1172 | 0 |
| N2 | 100 | 0.27 | 2.13 | 1.06 | 100 | 100 | 83 |
| SO2 | 1180 | 6.65 | 374.21 | 229.39 | 1180 | 0 | 0 |
| ClH | 33 | 0.27 | 7.05 | 2.65 | 33 | 33 | 20 |
| H2S | 212 | 0.48 | 135.60 | 26.96 | 212 | 212 | 94 |
| None | 208 | 0.53 | 630.70 | 21.64 | 208 | 53 | 7 |
| Total | 10,000 | 10,000 | 5572 | 4206 | |||
Note. — Stable condensates are defined by an activity of unity.
Figure 5 shows the atmospheres at 280 K that were previously in equilibrium with a fully molten mantle and summary statistics are presented in Table 3. The presence of condensates buffers the atmospheric composition and produces well-defined linear trends and clusters in the data. Therefore, and unlike at high temperatures, buffering of gas species fugacities by condensates means that atmospheres at 280 K can be grouped into discontinuous ”families,” of which we identify four types (see plots of log10SO2 VMR in Figure 5). Among these families, CO2-rich atmospheres are dominant (83% of the total) as revealed by the sharp peak in the frequency distribution at g mol-1. These develop from cooling of high temperature atmospheres produced across most of the initial O2 range, except in the most reduced (IW = -3) and oxidized (IW = +3) cases. Initially reduced atmospheres result in a subsidiary peak associated with CH4(g)-rich types (5% of the total) and a tail toward H2(g)-rich atmospheres (1.5% of the total). Another abundant gas in these atmospheres is NH3(g). In general, however, there is a notable dearth of atmospheres with molar masses below about 40 g mol-1 due to the (1) high C/H starting state caused by H2O(g) dissolution in the magma ocean, and (2) the stability of CO2(g) relative to CO(g) at low temperatures. Finally, oxidized starting conditions (i.e., IW = +3 to +4, see Section 3.1.2) extend the tail of the distribution to high molar masses, defining the fourth family associated with CO2–SO2-bearing atmospheres, which represent 8% of all cases.
These atmospheres can also produce up to 0.3 Earth-ocean equivalents of -sulfur (Figure 5). Crucially, the starting conditions that favor large amounts of sulfur precipitation (oxidized) are different from the conditions that favor significant water, graphite, and ammonium chloride precipitation (IW and below). This arises because S is more abundant in high-temperature atmospheres at oxidizing conditions, while C and particularly H are more abundant at more reducing conditions (see Section 3.1.2, Figure 4). There is also a trade-off between graphite and water, since high C/H in the starting conditions promotes larger graphite inventories, whereas low C/H forms larger water oceans. In the family of atmospheres with the lowest molar masses () produced by initially low C/H ratios and high H mass fractions, graphite saturation is not reached. The median total atmospheric pressure is roughly 65 bar and the maximum is a few kbar for oxidized starting conditions with high C/H and high initial number of water oceans.
Consideration of cases in which the mantle is only 10% molten leads to cooled atmospheres that have greater mixing ratios of H-bearing species. This leads to a more continuous distribution in across the simulation space (Figure B4), in which the heights of peaks associated with H-bearing species (notably H2(g) and CH4(g), given that H2O is always condensed) compared to that of CO2(g) are relatively greater than for the 100% molten cases. Nevertheless, the four families of atmospheres previously identified for the fully molten case are again evident, albeit slightly less distinct. CO2 atmospheres are similarly dominant, with a tail extending to SO2 atmospheres. Table 3 also reveals the formation of H2S-rich atmospheres for the molten case, although they are low in number (2% of the total).
3.2 K2-18b and similar sub-Neptune-sized planets
3.2.1 Context and parameters
Many sub-Neptunes likely harbor massive magma oceans beneath large insulating atmospheres, a configuration that corresponds to so-called ”gas dwarfs.” Such a structure has been proposed for the canonical sub-Neptune K2-18b (Shorttle et al., 2024; Wogan et al., 2024), although an alternative hypothesis suggests it might support clement conditions at the base of its atmosphere (Madhusudhan et al., 2023). Other sub-Neptune-sized exoplanets are currently being observed with JWST (e.g., TOI-270 d, GJ 1214 b), and preliminary observations reveal absorption features of various gas species . However, whether or not these planets possess a magma ocean surface, a temperate surface that could support liquid water, or no defined surface at all (e.g., Young et al., 2024b), remains unknown. Distinguishing between these scenarios requires self-consistent modeling frameworks that adequately capture the governing physics and chemistry of both the interior and the atmosphere. To this end, Atmodeller allows us to test the gas dwarf model with an important addition that has often been neglected in previous work on sub-Neptunes (Schlichting & Young, 2022; Misener et al., 2023; Young et al., 2023; Shorttle et al., 2024; Rogers et al., 2024; Rigby et al., 2024), though see Tian & Heng (2024); Seo et al. (2024): the role of nonideality, in which the volatile species at the conditions of the magma ocean–atmosphere interface do not obey the ideal gas law. Rather, real gases decouple their partial pressure (which is related to molar abundance, mass balance, and mechanical pressure) from their thermodynamic pressure, which regulates equilibrium chemistry and solubility (Section 2.3).
We simulate K2-18 b as a gas dwarf and demonstrate the effects of nonideality on sub-Neptune atmospheres. K2-18b has a total mass of kg (8.63 ) and an estimated mantle mass of kg (assuming the same core/mantle mass ratio as the Earth) (Benneke et al., 2019). K2-18b’s surface radius, assuming it has an optically thick atmosphere overlying a magma ocean surface, is estimated to be 11225 km (1.76 ; assuming an Earth-like interior composition and the super-Earth mass–radius relation taken from Table 2 of Hakim et al., 2018). The surface temperature at the magma ocean–atmosphere interface is set at 3000 K, We further note that this temperature is at or above the upper limit of the calibration for current solubility laws and real gas EOS. The mantle melt fraction is set to 100%. We consider six gas species in the atmosphere, each with their corresponding solubility laws, necessarily assuming the mantle composition is basaltic to remain internally consistent: H2O (Dixon et al., 1995), H2 (Hirschmann et al., 2012), CO (Yoshioka et al., 2019), CO2 (Dixon et al., 1995), and CH4 (Ardia et al., 2013). O2 is also included as a gas species. For the cases with real gases, the CORK model is used for H2O, an ab initio EOS model for H2 (Chabrier & Debras, 2021), and a corresponding states model for CO, CO2, and CH4 (Shi & Saxena, 1992) (Table 2). O2 is assumed to behave as an ideal gas in all cases, which does not introduce any uncertainty owing to its vanishingly low fugacities under the range of conditions studied.
To explore the effects of nonideality, we perform two series of simulations in which volatiles can partition between the silicate mantle and the atmosphere. The total hydrogen inventory is varied from 0.1% to 3% of K2-18b’s mass (Figures 6 and 7) assuming both real and ideal gas behavior. In these simulations, all other parameters are held constant: K, C/H (by mass) = 0.32 ( the solar ratio, as recently inferred for K2-18b, Wogan et al. (2024)) and O2 at IW = -3. The equivalent cases, in which each of the other two variables (IW and C/H ratio) are varied while the others are held constant, are shown in Figures G1 and G2, respectively. All simulation data are available to download (Bower, 2025).
3.2.2 Volatile storage in the mantle
The key result is that for planets with large volatile inventories, the mass fraction of volatile elements (H+C) retained in the envelope (up to 4% of the planet mass initially added, ) decreases as solubility increases, and decreases even further when real gas behavior is taken into account. In Figure 6a, a case without solubility (not shown) would produce a 1:1 relationship between the envelope volatile mass fraction (-axis) and the total volatile mass fraction (-axis). When ideal gas behavior and solubility are considered (gray dashed line), the relationship is approximately 2:3, and when both real gas behavior and solubility are included (gray solid line), the relationship is closer to 1:3. The ideal with solubility case overestimates the volatile mass fraction in the envelope (Figure 6a), and hence the total atmospheric pressure by a factor of 2 compared to the real with solubility case (Figure 7a). Relative to a case without solubility, the envelope volatile mass fraction is reduced by 2/3 when considering both H- and C-bearing species solubilities and real gases. For cases that do consider solubility, assuming a total hydrogen mass equivalent to 3% of K2-18b’s mass results in a 50% decrease of the total atmosphere pressure for the real case (12 GPa) compared to the ideal case (27 GPa). Therefore, these cases demonstrate that both solubility and nonideality have comparable influences in determining the first-order characteristics of sub-Neptune atmospheres.
The mass fraction of H+C in the envelope predicted by the ideal and real cases with solubility begins to diverge above 0.2–0.3 total volatile mass fraction (compare gray dashed and solid lines, Figure 6). This is because, as the hydrogen mass fraction of the planet increases, the atmospheric surface pressure increases, causing the gas fugacity coefficients, , to deviate more strongly from ideality (Figure 7d). Because the temperatures and pressures far exceed the critical points of the gas species considered here, repulsive forces between the molecules predominate, and volumes (), and hence exceed those of ideal gases. For a constant mass budget of each element, this property increases the fugacity, and hence solubility in silicate liquids, since solubility is proportional to thermodynamic pressure (i.e., fugacity) not partial pressure (Figure 7c). Consequently, the enhanced solubility of volatile elements, primarily hydrogen species (H2 and H2O), limits the amount of H and O (and to a lesser extent C) in the atmosphere and acts to buffer increases in the total atmospheric pressure (Figure 7c,d).
At fixed O2(IW = -3) and C/H (0.32 by mass; 100 solar), for low H mass fractions, CO(g) is the dominant species, but is exceeded by H2(g) across all hydrogen mass fractions 0.2% in the ideal case, whereas CH4 exceeds H2 above 0.3% H mass fraction in the real case. The predominance of CH4(g) results from the interplay between total pressure and the fugacity coefficients of the major atmospheric species, as described by the homogeneous equilibrium:
where, at equilibrium, and at constant O2:
Therefore, CH4 is proportional to , whereas scales with . Based on this reasoning, at constant bulk composition, temperature and O2, CH4(g) should become dominant as increases. However, this is not observed in the ideal case, where H2 exceeds CH4 across the entire pressure range (Figure 7a). In ideal cases, H2 remains stable because the addition of gas with C/H (by mass) = 0.32 gives rise to C/H ratios in the atmosphere that are still too low (i.e., 1 by mass; Figure 7b) to form significant quantities of methane (Figure 7a). On the other hand, for real cases, the increase in H2 results in its increased dissolution relative to the ideal case (Figure 7c) while CH4 never dissolves to any substantial extent, causing a commensurate increase in C/H of the atmosphere ( 2.5 by mass, Figure 7b) that is closer to the stoichiometry of CH4 (C/H = 3 by mass). Although CH4 is in the range 103 and H2 20–30, since H2 depends upon H and CH4, the effect of relative differences in essentially cancel one another out.
The aforementioned calculations were performed for a fixed planetary mass and radius. However, these quantities are only determined within limits for sub-Neptunes (at best 15% relative for mass and 3.5% for radius, Luque et al., 2022). Hence, we now determine, for a fixed hydrogen mass fraction, C/H, O2 and surface temperature, how variations in a planet’s surface radius or mass can influence the atmospheric speciation we infer when employing real gas EOS. As shown in Figure 8a, in the real case, for a fixed mass of 8.63 , the atmosphere will be dominated by CH4 below 2 and by H2 above 2 . This contrasts with the ideal scenario, in which H2 is always the dominant species. Similarly, at constant radius (1.76 R⊕), the atmosphere is H2 dominated below 7 M⊕ and CH4 dominated above 7 M⊕ for real gases (Figure 8b). Hence, typical uncertainties in mass and radius (Figure 8, gray horizontal bars) can bracket the transition regime between H2 and CH4 dominated atmospheres, highlighting the importance of improving mass and radius measurement uncertainties to accurately interpret atmospheric observations. Mass and radius both influence the atmospheric pressure at the surface, which in turn impact fugacity coefficients, through the relationship:
where is the total mass of the atmosphere, is the acceleration due to gravity, and is the surface radius. Because pressure is proportional to 1/, even small changes in surface radius have tangible effects on total pressure, for a constant atmospheric mass budget. Conversely, the linear dependence of on (and hence on the planetary mass, ) results in a more modest mutual dependence. , and increasing pressure for a constant total mass budget results in higher dissolution (since , where is any given gas species), giving rise to the monotonic decrease in H2 (and hence total ) with increasing or with decreasing in the ideal case. Conversely, increasing leads to increases in (analogous to the mechanism discussed with respect to Equation 6, above), promoting additional H dissolution as molecular H2. The net result is sub-Neptunes are expected to be CH4-rich for cases in which mean densities are roughly 6500 kg m-3 at IW = -3, H=3% , and bulk C/H=100 solar. Figures G1 and G2 illustrate the effects of varying oxygen fugacity from -6 ¡ IW ¡ 0 (at constant C/H = 0.32 and H mass fraction = 1% ) and total C/H mass ratio (at constant O2 = IW = -3 and H mass fraction = 1% ), respectively, on volatile speciation in the atmosphere and dissolution in the interior assuming real and ideal gas behavior, for K2-18b-like cases.
4 Discussion
4.1 Earth-sized planets
For Earth-sized planets, our results show that the dominant atmosphere at temperate conditions that emerges after planet formation is CO2-rich with a total surface pressure around 100 bar. Such an atmosphere forming upon cooling on the early Earth from a magma ocean state was first hypothesized by Holland (1984), computed by Zahnle et al. (2007), and demonstrated by Sossi et al. (2020) for a magma ocean initially equilibrated at its estimated O2 close to the IW buffer. CO2-rich atmospheres are compatible with that of Venus at the present day, both in composition and total pressure. The composition is also similar to that of modern Mars, hinting that this mode of atmosphere formation may have been commonplace on rocky planets in our solar system. That the Earth does not currently possess a 97:3 CO2–N2 atmosphere points to the operation of geochemical cycles that drew down carbon from the atmosphere through weathering by the Urey reaction to form carbonates (Sleep & Zahnle, 2001; Zahnle et al., 2007) that are subsequently recycled into the interior, thus evolving its atmosphere away from CO2-rich over geological timescales. This is because carbon is abundant, CO2 is strongly favored at 280 K, and water condensation buffers the emergence of hydrogen-dominated atmospheres even when total C/H is low (as with atmospheres derived from a partially molten starting state, Appendix B).
Condensates at planetary surfaces, such as water, graphite, -sulfur, and ammonium chloride, are of primary interest for the establishment of geochemical cycles (Hirschmann, 2018; Dasgupta et al., 2024) and a habitable environment (Benner et al., 2020; Mrnjavac et al., 2023; Rimmer & Shorttle, 2024). Figures 9 and B5 show atmospheres at 280 K above surfaces of an Earth-like planet together with stable condensates with which they are in equilibrium, derived from a fully molten and partially molten mantle, respectively. Each panel mandates that a particular condensate is stable with a minimum mass (shown on the -axis), where the upper row reveals the number of models associated with a different dominant (VMR 50%) species as a function of the number of Earth-ocean equivalents by mass (EO) condensed. The lower row shows the variety of atmospheric species for dominant classes of atmospheres that support condensates, where either CO2 or CH4 is chosen as the dominant species (Table 3). The conditions for water oceans and graphite to be stable are comparable, with two classes of atmospheres emerging as determined by a VMR greater than 50%: CO2-rich and CH4-rich, with the former representing 83% of all atmospheres, whereas the latter only around 5% (e.g., Figure 9). For both classes, the atmospheres have median total surface pressures of a few tens of bar, and due to buffering imposed by C(cr) and H2O(l) condensates, the partial pressures of other C- and H-bearing species are relatively constant, regardless of the condensate reservoir mass.
Small amounts of water can be stable beneath atmospheres dominated by SO2, N2, H2, and ClH, although these are all less prevalent than CO2 and CH4 (Figure 9a). Conditions that permit the condensation of solid ammonium chloride, ClH4N(cr) mirror those for water and graphite, reflecting the greater availability of NH3 and HCl in reduced atmospheres relative to their oxidized counterparts (Figure 2). Ammonium chloride most commonly condenses in small mass fractions (0.0025 EO) beneath CO2-rich atmospheres, but the highest ClH4N(cr) mass fractions (0.006 up to 0.012 EO) occur, albeit rarely (1% of all models), in equilibrium with H2. Conversely, -sulfur stability occurs almost exclusively beneath CO2-rich atmospheres, in which the partial pressure of CO2 is on average higher than in equilibrium with water and graphite (compare the orange lines in Figure 9e and g). The CO2-rich atmospheres in equilibrium with -sulfur also frequently have several tens of bar of SO2 and only 10-11 bar of H2S. Note that SO2(g) is almost never the major gas phase above -sulfur because most of the S budget is hosted in the condensed phase. In total, water, graphite, -sulfur, and ammonium chloride coexist in around 40% of the models, albeit at potentially low abundance. Nevertheless, our results show that many early atmospheres are conducive to the formation of condensates on temperate planetary surfaces (here 280 K), should these atmospheres be able to cool isochemically and reach chemical equilibrium with stable condensates.
Figure 10 summarizes the atmospheric abundances (ppm) above planetary surfaces that simultaneously support water, graphite, -sulfur, and ammonium chloride. The objective is to identify characteristic atmospheric signatures that may be used to identify planetary surfaces with a higher propensity for habitability (Krissansen-Totton et al., 2018; Schwieterman et al., 2019; Konrad et al., 2024; Tokadjian et al., 2024; Borges et al., 2024; Young et al., 2024a). To satisfy joint condensate stability, reduced species, such as H2, NH3 and CH4 are low in abundance—almost always below the ppm level—because otherwise -sulfur does not form. This comes from the fact that these species are present in the atmosphere only at very low O2 at the high-temperature initial conditions, at which S is highly soluble. Therefore, S is never present in sufficient amounts in these atmospheres to be able to condense upon cooling to a low temperature. Conversely, oxidized initial conditions that produce significant SO2 lead to final, elevated O2s of these atmospheres that prevent the condensation of graphite via CO2(g) = C(cr) + O2(g), even at high CO2. Consequently, any detection of SO2 in an atmosphere is incompatible with the presence of graphite. On the other hand, SO2 is highly unlikely to be the dominant atmospheric gas in the presence of H2O(l) or -S(cr). Species whose fugacities are less redox sensitive cannot readily be used to discriminate between stable versus unstable condensates (e.g., H2O and N2). Figure B6 illustrates the same analysis but excludes -sulfur and ammonium chloride, and hence only considers the joint stability of water and graphite. In this scenario, reduced species can be accommodated at higher abundances (e.g. H2, NH3), and notably CH4-rich atmospheres are compatible with both water oceans and graphite at planetary surfaces.
Our analysis provides a geological baseline for the expected states of planetary atmospheres after they emerge from the magma ocean stage, postformation but prior to modification by geochemical cycling, atmospheric (photo)chemistry and escape, and later, possibly life. It should be noted that these computations assume chemical equilibrium is reached at 280 K in the absence of chemical modification of the atmosphere following the magma ocean stage. In practice, the timescales over which the atmospheres cool may be sufficiently short so as to permit chemical kinetics and photochemical processes (Arney et al., 2016; Zahnle et al., 2020; Drant et al., 2025), in addition to continued volcanic outgassing (Ortenzi et al., 2020; Guimond et al., 2021; Liggins et al., 2022) and/or atmospheric escape/erosion (e.g., Krissansen-Totton et al., 2024) to invalidate our dual assumptions of chemical equilibrium at a constant bulk composition. Hence, these results should be regarded as an end-member case, or to provide the initial atmospheric composition to inform the subsequent time evolution of cooling. We anticipate that our analysis can inform a prior expectation in modeling studies and statistical analyses, in which observations can be used to quantify whether a planetary atmosphere has evolved away from its earliest outgassed state.
The establishment of geological cycling may require an early reservoir of water at the surface or in the upper mantle (e.g., Tikoo & Elkins-Tanton, 2017) to alter the rheology and partial melting of rocks (Korenaga, 2020). In this regard, surfaces that emerge postformation with condensates, notably liquid water, may already be primed to cycle volatiles with the interior and thereby modify the size and composition of the atmosphere. Conversely, for surfaces that emerge without condensates, there could be a higher propensity to remain locked in an inactive state since geochemical cycling cannot initiate, at least not immediately. Table 3 reveals that most (91%) Earth-like planets that emerge from the magma ocean stage have the prerequisite conditions for surficial water, with at least 56% of planets also accommodating graphite. Since condensates buffer the atmospheric speciation (due to the equivalence of chemical potentials, an intensive variable), finding atmospheres that are compatible with stable condensates will not directly allow the size of those condensate reservoirs to be determined.
4.2 Sub-Neptunes
Our analysis permits new insights on two key aspects of the nature of sub-Neptunes: (1) the influence of atmospheres on their mass-radius relationships and demographics (e.g., Bean et al., 2021; Luque & Pallé, 2022) and (2) the identity and abundance of gas species in equilibrium with magma oceans, providing grounds for comparison with spectroscopic observational constraints (e.g., Madhusudhan et al., 2023; Benneke et al., 2024).
4.2.1 Nonideality and the ”Fugacity Crisis”
Regarding point 1, Kite et al. (2019) proposed that a ”fugacity crisis,” in which the enhanced dissolution of H2 at high surface pressures (1–8 GPa) hinders atmospheric growth, can explain the drop-off in exoplanet abundance above 3 observed in exoplanet demographics studies (i.e., the ”radius cliff”) (Fulton & Petigura, 2018); though see also Schlichting & Young (2022) for a different interpretation. Our work goes beyond the H2 nonideality case considered in Kite et al. (2019), by incorporating additional volatile species (e.g., H2O, CH4, CO, and CO2) along with their dissolution into the interior and real gas EOS. Despite the relatively large mass of envelopes (2%–3% of the planetary mass) required to explain the mass-radius relationships of sub-Neptune exoplanets in the absence of interior dissolution, these quantities are modest, on a molar basis, with respect to the proportion of condensed mass in the planet. As such, in the end-member case of a solar gas interacting with a rocky sub-Neptune interior, the equilibrium H/O and C/O ratios will be modified from their solar values (2000 and 0.55 by moles, respectively, Asplund et al., 2009) via the mutual solubility of O in the envelope and C and H in the mantle. Since the H/O and C/O ratios of nominally rocky material are 1 (H/O and C/O ratios in the BSE are 0.005 and 3.6 10-4 by moles, respectively, Palme & O’Neill, 2014), chemical equilibrium will result in a decrease in H/O and C/O in the envelope relative to solar values (see also Young et al., 2023).
Here, because only volatile element mass balance is considered (i.e., there is no notion of the chemical composition of the condensed phase other than for C, H, O, and N), this process is simulated by considering a higher O2 (IW = -3) in our sub-Neptune models than that imposed by the solar nebula gas (IW = -6.5, Grossman et al., 2008). Should O2 at the magma ocean–envelope interface fall below IW = -3.5, then H2 predominates in the gas phase (envelope), even for C/H ratios 100 solar, though CH4 remains a minor species (Figure G1). As noted by Kite et al. (2019); Schlichting & Young (2022), however, the experiments underpinning the solubility law for H2 dissolution in silicate liquids were performed over a limited temperature (1673–1773 K) and pressure (up to 3 GPa) range (Hirschmann et al., 2012); the temperature range is notably much lower than the 3000 K assumed for equilibrium at the interface. Therefore, how the equilibrium constant of the reaction H2(g) = H2(l) depends on temperature, and pressures above 3 GPa, is unconstrained, yet is essential to assess the ”fugacity crisis” hypothesis.
4.2.2 Role of Nonideality in Inferring Atmospheric Chemistry
A key feature of spectroscopic determinations of the speciation of gases in the atmospheres of temperature sub-Neptunes with equilibrium temperatures in the range 250–400 K, such as K2-18b ( = 255 K) and TOI-270d ( = 354 K), is the ubiquity of CO2(g) and CH4(g) (Madhusudhan et al., 2023; Benneke et al., 2024). Mixing ratios of these species are of the order 1–2% each, and their ratio, is roughly unity (Madhusudhan et al., 2023; Benneke et al., 2024), with evidence for H2O(g) at similar levels for TOI-270d but unresolved (below 10-3) in K2-18b. In all cases, the background gas is inferred to be H2. Although these transmission spectroscopy measurements probe the upper atmosphere (near 10-2–10-4 bar, Benneke et al., 2024; Madhusudhan et al., 2025) and the speciation cannot therefore be directly compared with that predicted in our models, they provide first-order constraints on the chemical degrees of freedom permitted in sub-Neptune atmospheres.
Inferred metallicities in the atmosphere required to fit the observed CH4 and CO2 abundances are of the order 100–200 solar (Wogan et al., 2024; Yang & Hu, 2024), which are consistent with 100 solar composition used in our simulations (Section 3.2). At these metallicities, we highlight that CH4(g) is expected to be a major species. Consequently, the C/H of the bulk atmosphere is expected to be 2.5 by mass (Figure 7), and its variation is antithetic with that of C/O, which increases monotonically from near zero to 300 as the fraction of H present as a function of planetary mass increases from zero to 3% (at constant IW = -3). Inferred C/H and C/O ratios (by mass) from the measurements of Madhusudhan et al. (2023); Benneke et al. (2024) are of the order 0.1–0.3 (30–100 solar) and 0.02–0.5, respectively. Hence, both are lower than those predicted in our simulations for a real gas in equilibrium with a magma ocean (Figure 7). Partitioning of carbon into a metal phase, which we do explicitly model, could reduce the atmospheric C/O ratio (Werlen et al., 2025).
The lower C/O ratios observed than predicted by our models leads to another discrepancy—the absence of CO2(g) in our simulations. As highlighted for terrestrial planetary atmospheres (Section 3.1), however, any CO(g) that is stable at high temperatures (here, 3000 K) is partitioned between CO2(g), C(cr). and/or CH4(g) at low temperatures (300–400 K, the temperatures in the photosphere recovered from transmission spectra) depending on the prevailing gas chemistry and total pressure. This implies substantial amounts of CO2(g) could form in the upper atmosphere upon cooling. Therefore, future work should concentrate on how gas-condensed phase equilibria act to modify the speciation of sub-Neptune atmospheres throughout the atmospheric column.
A simplified version of this exercise was carried out by Yang & Hu (2024), in which speciation at the pressures of observation in the atmospheres of K2-18b and TOI-270d were extrapolated to 200 bar. In order to match the CH4/CO2 proportions in the upper atmosphere, these authors inferred H2/H2O ratios at 200 bar to be roughly unity for K2-18b and 0.5 for TOI-270d. Assuming these ratios are representative of those at any magma ocean/atmosphere interface, this would correspond to IW in the range 0 to 1. These O2s are similar to, or more oxidized than, those inferred for small, differentiated rocky bodies in the solar system (Wadhwa, 2008). Such high oxygen fugacities would seem at odds with the substantial quantities (1%–3% by mass) of solar-like gas inferred to have been accreted by sub-Neptunes. Indeed, we predict C/H ratios that are too high (2.5) at these O2s, and H2 would no longer be a predominant constituent of the atmosphere (Figure G1). Therefore, preliminary indications suggest that either (1) bulk planetary C/H ratios should be relatively low (10-1 by mass), either by processes that fractionate C from other atmosphere-forming elements during accretion relative to solar or its preferential sequestration into sub-Neptune cores, and/or (2) the lower atmosphere has a distinct chemical composition with higher C/H and C/O than does the upper atmosphere sensed by transmission measurements (e.g., by inefficient mixing or condensation of graphite in the atmospheric column), in order to maintain simultaneously low C/H (0.1–0.3) and C/O (0.02–0.5) ratios observed in temperate sub-Neptunes.
5 Conclusions
For Earth-sized rocky planets such as TRAPPIST-1e, the atmosphere during the magma ocean stage is likely CO-rich due to two key factors. First, CO(g) is thermodynamically favored at 1800 K across a broad range of oxygen fugacities. Second, carbon-bearing species are significantly less soluble than hydrogen-bearing species—particularly H2O(g)—,leading to an elevated atmospheric C/H, often exceeding the bulk C/H by several orders of magnitude. As a result, hydrogen-bearing species are depleted in the atmosphere, even when the mantle is only partially molten. In a fully molten mantle, the further depletion of hydrogen by the sequestration of H2O(g) into the liquid leads to atmospheres that are defined by entirely continuous mixing ratios of their constituent species, favoring atmospheres dominated by (1) heavy carbon species (predominantly CO(g)), (2) CO2(g) SO2(g)-bearing varieties at high O2s (IW = +3 to +4), or (3) H2(g) CH4(g)-bearing types at low O2 (below IW = -3).
Based on the equilibrium results of atmospheres with molten mantles, we cool the atmospheres to 280 K and permit (if thermodynamically stable) the formation of water (oceans), graphite, -sulfur, and ammonium chloride. Unlike at high temperatures, the presence of condensates buffers gas mixing ratios to produce discrete ”families” of atmospheres. Notably, CO2-rich atmospheres are not only widespread (83% of cases) but also uniquely capable of supporting a wider variety of surface condensates compared to those dominated by other species. A small fraction (5%) are CH4-rich atmospheres that can sustain both water and graphite, while roughly 8% are CO2–SO2-rich in equilibrium with native sulfur. The remainder (1%) are H2-rich. Depending on the conditions during early equilibration with a magma ocean, atmospheres rich in N2, HCl, or H2S can also host diverse surface condensates; however, such cases comprise only 1% of all cases. In short, water condensation buffers the partial pressures of hydrogen-bearing species in the atmosphere, such that even low C/H does not produce substantial hydrogen atmospheres except at the most reduced conditions where H2 can dominate. Thus, for both high and low temperature atmospheres, carbon always wins.
Our simulations can be considered as geological reference cases—the expected state of a rocky planet after its earliest evolution. From this state, subsequent processes such as geological outgassing and atmospheric (photo)chemistry, dynamics, and escape can modify the atmosphere over geological timescales. Hence, our results can be used to inform the priors and initial conditions of other models, as well as provide an observational baseline for rocky planets to determine if they have evolved away from their birth state. In this regard, the simulations can inform observational studies and instrument requirements to investigate the atmospheres of temperate rocky planets.
For sub-Neptune-sized planets, we performed nonideal gas-phase chemistry simulations including dominant C- and H-bearing species at conditions relevant to the magma ocean–envelope interface. The combination of solubility and nonideality conspire to stabilize CH4 as a major gas species for a range of parameter choices that are reasonable given current observational constraints for sub-Neptunes. As for Earth-sized rocky planets, this result is a natural outcome of the high solubility of hydrogen-bearing species (notably molecular hydrogen for sub-Neptunes) compared to carbon-bearing species, a result that is exacerbated due to nonideality at the high-temperature (3000 K) and pressure (several GPa) conditions assumed to occur at the interface. Furthermore, the transition between a H2 and CH4 atmosphere can occur within the range of observational uncertainties in mass and radius, even for a fixed volatile budget and O2, thereby providing a new challenge to infer atmospheric composition. Given detections of carbon-bearing species in the atmospheres of sub-Neptunes, such as K2-18b and TOI-270d, our findings emphasize the need to include and realistically model gas envelopes throughout their extent.
To achieve these results, we devised an open-source code Atmodeller, built on JAX, to probe the nature of atmospheres above rocky planets from Earth size to sub-Neptune size. The code includes solubility laws, real gas EOS, and thermodynamic data. Leveraging the functionality of the JAX Python library allows the code to be both user friendly and highly performant, and community development of the code is encouraged.
acknowledgments
We thank Daniel Kitzmann, Patrick Kidger, and Johanna Haffner for discussions. D.J.B., M.A.T., and P.A.S. were supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract No. MB22.00033, a SERI-funded ERC Starting grant “2ATMO.” P.A.S. also thanks the Swiss National Science Foundation (SNSF) through an Eccellenza Professorship (#203668). M.A.T. is supported by NASA through the NASA Hubble Fellowship grant #HST-HF2-51545 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. K.H. acknowledges the FED-tWIN research program STELLA (Prf-2021-022) funded by the Belgian Science Policy Office (BELSPO) and the FWO research grant G014425N. M.T. acknowledges financial support from the Chair of Theoretical Astrophysics of Extrasolar Planets at LMU Munich.
Appendix A Sampling Strategy and Parameter Choices
The logarithmic sampling strategy adopted to explore the atmospheric diversity of TRAPPIST-1e is physically motivated as follows. Oxygen fugacity (O2) exerts a fundamental control on atmospheric composition, particularly through its influence on redox speciation; for example, determining the relative abundances of reduced versus oxidized species such as H2 and H2O. Thermodynamically, O2 is defined by an equilibrium reaction, and when referenced to Fe–FeO (IW) equilibrium (Equation E1) is conventionally expressed as a offset relative to the IW buffer. This convention naturally supports the use of a log-uniform prior, especially within a thermochemical framework where partial pressures, fugacities, and equilibrium constants combine multiplicatively (and thus additively in log-space). Observationally, the present-day O2s of rocky bodies in the solar system spans several orders of magnitude from roughly 5 units below IW (Mercury) to 2 units above IW (carbonaceous chondrites) (e.g., Grewal et al., 2024), with several extrasolar rocky bodies showing comparable redox states (Doyle et al., 2019). Early Earth conditions during the magma ocean stage were close to IW (Sossi et al., 2020).
Therefore, to capture this diversity, we center our log-uniform prior on zero offset relative to IW (anchored to early Earth) and allow sampling across a broad range of units. While no prior is without limitations, we adopt this scale-invariant formulation because it more faithfully represents ignorance over the parameter’s order of magnitude. It offers the dual advantages of anchoring to a well-motivated reference point (early Earth) while permitting exploration beyond the present-day solar system constraints on O2, which is warranted by the potentially greater diversity of exoplanetary environments. For consistency with the log-uniform prior, we apply the same sampling strategy to the other two parameters: the total hydrogen inventory and the C/H mass ratio are each varied across one order of magnitude on either side of an Earth-like reference value. The upper bound of the total hydrogen inventory is motivated by planet formation simulations, which show that terrestrial planets can accrete several times Earth’s water content (Raymond et al., 2007). The lower bound, by contrast, reflects scenarios of limited delivery or substantial early loss.
The corner plots display bivariate scatter plots (e.g., Figure 1), which directly reflect the model outcomes and are therefore insensitive to the choice of prior, depending only on the sampled parameter ranges. In contrast, the likelihood contours, marginal distributions, and derived summary statistics (e.g., Table 3) are prior dependent. A uniform distribution might appear to be a simple alternative prior, but its inadequacy for O2 is evident: when sampled uniformly between and , the mean value in space is 4.7, which is heavily skewed toward the upper bound and far outside the range characteristic of the early Earth or any known rocky body in the solar system. New observations and models may eventually reveal the true distributions of the parameters that govern atmospheric diversity, but, in the absence of such knowledge, a logarithmic sampling strategy remains the most reasonable and defensible choice.
Appendix B Partially molten mantle
We repeat the analysis presented in Section 3.1 for a partially molten TRAPPIST-1e with a mantle melt fraction of 10%, whereby the influence of solubility is lessened compared to a fully molten mantle (compare Figures 1, 2, 4, 5 to Figures B1, B2, B3, B4, respectively). Even with just 10% melt fraction, C/H in the atmosphere is boosted and its distribution broadened compared to total C/H (Figure 3). Compared to a fully molten mantle, a partially molten mantle permits a greater diversity of atmospheres since a broader range of gas mixtures comprised of hydrogen and carbon species are possible. Although larger H2O atmospheres can occur, the median H2O in the atmosphere is still only 11 mol% versus 1 mol% for a fully molten mantle. The mean molar mass (Figure B1, bottom right) reflects this greater atmospheric diversity, where a subsidiary peak around 5 g mol-1 represents H2-dominated atmospheres at reduced conditions in a mixture with CH4 and CO. The CO peak at 28 g mol-1 is evident, albeit now supplemented by mixtures of H2O (18 g mol-1) and CO2 (44 g mol-1) in approximately equal proportion. The molar mass distribution also has an extended tail, driven by the prevalence of CO across a broad range of oxygen fugacity as well as the emergence of CO2 (44 g mol-1) and SO2 (64 g mol-1) as prevalent oxidized species. Regardless, the thermodynamic stability of CO over a range of O2 ensures that it remains a dominant species.
For a partially molten mantle, dissolved H is more than 60% for models above the IW buffer due to abundant H2O which is soluble. For reduced conditions below the IW buffer, H2 is more abundant than H2O, and, since H2 is not as soluble as H2O, less total hydrogen is dissolved, between about 10% and 60%. For comparison, dissolved C never exceeds around 3% regardless of whether the reduced (CO) or oxidized (CO2) species is most abundant. We recall that total C/H is an input parameter, which is varied between -1 and 1 log10 units by mass, which corresponds approximately to between -2 and 0 log10 units by moles. This gives rise to a median log10 C/H in the atmosphere of -0.2 and C/H can reach several hundred for the most oxidized conditions. Hence even with a small mantle melt fraction (10%) total C/H can differ from atmospheric C/H by around two orders of magnitude.
Figure B4 shows the atmospheres at 280 K derived from the partially molten starting state. Compared to those derived from the fully molten starting state, the atmospheres exhibit the same complex relationships, albeit the solution space is enlarged for certain quantities. This is because the smaller initial C/H allows H2 and CH4 atmospheres to be more prevalent at reduced starting conditions which can produce mixtures of hydrogen and carbon species. Nevertheless, CO2 atmospheres are dominant in similarity with atmospheres derived from the fully molten starting state. H2O is only present in the atmosphere below the percent level since water condenses to produce a median surficial reservoir of 0.4 Earth oceans by mass (EO) with a maximum inventory around 5.5 EO, which is almost a tenfold increase compared to the fully molten starting condition (Figure 5). -sulfur condensation can produce up to around 0.4 EO, although an expected amount is closer to 0.1 EO. In contrast, the total graphite reservoir is usually around 0.3 EO, but for reduced starting conditions can reach up to 10 EO. Large reservoirs of all three condensates can coexist, usually for starting conditions around the IW buffer.
The overall trends of atmospheres above planetary surfaces with stable condensates are comparable to the fully molten starting state in terms of the dominance of CO2 and CH4 atmospheres (compare Figures B5 and 9). However, there are a few notable differences. First, most water oceans are stable beneath CO2 atmospheres and most graphite reservoirs are stable beneath CH4 atmospheres. This trend is reversed compared to the atmospheres derived from a fully molten starting state. Second, H2S atmospheres are absent for the fully molten starting state, but now feature above all three condensate reservoirs. Third, SO2 atmospheres can support up to 3 EO of water as well as 0.15 EO of -sulfur. Finally, N2 atmospheres, although low in number, are also compatible with large reservoirs of water and -sulfur. Figure B6 shows the atmospheric speciation in equilibrium with stable reservoirs of water and graphite only, which excludes -sulfur and ammonium chloride compared to Figure 10 .
Tables B1 and B2 summarize the atmosphere speciation for TRAPPIST-1e at 280 K for atmospheres with stable water, graphite, and -sulfur, for fully molten and partially molten starting states, respectively (Figure 10).
| Species | All Condensates Stable (ppm) | At Least One Condensate Unstable (ppm) | ||||||||
| Min | Max | Q1 | Q2 | Q3 | Min | Max | Q1 | Q2 | Q3 | |
| H2O | 1.3e+00 | 1.8e+04 | 2.8e+01 | 8.6e+01 | 2.5e+02 | 1.1e+00 | 8.8e+04 | 6.5e+01 | 2.5e+02 | 9.4e+02 |
| H2 | 1.4e-08 | 5.8e-02 | 1.4e-06 | 7.8e-06 | 4.0e-05 | 3.4e-16 | 9.5e+05 | 1.6e-14 | 5.5e-05 | 8.5e-04 |
| CO2 | 1.6e+05 | 1.0e+06 | 9.7e+05 | 9.9e+05 | 1.0e+06 | 1.1e-30 | 1.0e+06 | 3.8e+05 | 8.1e+05 | 9.6e+05 |
| CO | 3.6e-08 | 2.7e-06 | 1.7e-07 | 3.0e-07 | 5.1e-07 | 6.2e-32 | 4.2e-06 | 4.4e-16 | 6.8e-10 | 9.3e-07 |
| CH4 | 7.5e-09 | 9.4e+00 | 3.7e-06 | 3.5e-05 | 3.1e-04 | 1.7e-43 | 1.0e+06 | 9.5e-40 | 4.3e-04 | 1.8e-02 |
| N2 | 1.8e-12 | 7.6e+05 | 3.1e+03 | 7.7e+03 | 2.6e+04 | 2.5e-46 | 8.4e+05 | 2.3e+03 | 1.4e+04 | 6.3e+04 |
| NH3 | 2.8e-13 | 1.7e-02 | 7.9e-08 | 5.6e-07 | 3.4e-06 | 1.4e-21 | 2.6e+04 | 7.2e-20 | 1.6e-12 | 1.5e-05 |
| H2S | 1.7e-02 | 7.1e+04 | 1.8e+00 | 9.6e+00 | 4.9e+01 | 4.2e-10 | 2.5e+05 | 1.9e-08 | 1.1e+01 | 1.2e+02 |
| SO2 | 3.8e-13 | 2.4e-12 | 2.3e-12 | 2.4e-12 | 2.4e-12 | 3.2e-45 | 9.7e+05 | 8.3e-14 | 1.4e-12 | 2.0e+05 |
| ClH | 3.8e-05 | 1.7e+05 | 8.3e-05 | 1.1e-04 | 1.5e-04 | 6.2e-14 | 9.3e+05 | 2.6e-04 | 2.3e+00 | 3.8e+01 |
Note. — Q1, Q2, and Q3 represent the first (25th percentile), second (median, 50th percentile), and third (75th percentile) quartiles, respectively. These data are represented in Figure 10(a,c), where interior dotted lines show Q1 and Q3 and dashed lines show Q2 (median).
| Species | All Condensates Stable (ppm) | At Least One Condensate Unstable (ppm) | ||||||||
| Min | Max | Q1 | Q2 | Q3 | Min | Max | Q1 | Q2 | Q3 | |
| H2O | 1.2e+00 | 3.7e+04 | 4.0e+01 | 1.7e+02 | 6.9e+02 | 1.0e+00 | 4.3e+04 | 3.5e+01 | 7.2e+01 | 3.0e+02 |
| H2 | 1.3e-08 | 7.9e-01 | 2.5e-06 | 2.2e-05 | 2.2e-04 | 2.1e-16 | 9.8e+05 | 5.0e-15 | 1.1e+00 | 6.5e+05 |
| CO2 | 2.3e-04 | 1.0e+06 | 9.3e+05 | 9.8e+05 | 1.0e+06 | 4.7e-32 | 1.0e+06 | 6.1e-26 | 2.7e-03 | 3.9e+05 |
| CO | 4.9e-12 | 3.9e-06 | 1.9e-07 | 3.9e-07 | 7.8e-07 | 4.8e-33 | 2.6e-06 | 3.0e-28 | 1.6e-16 | 3.0e-15 |
| CH4 | 6.8e-09 | 2.0e+05 | 7.7e-06 | 1.4e-04 | 3.1e-03 | 4.2e-43 | 9.9e+05 | 1.2e-40 | 5.6e+04 | 4.8e+05 |
| N2 | 5.2e-28 | 8.4e+05 | 7.4e-12 | 4.8e+03 | 3.8e+04 | 8.0e-49 | 8.8e+05 | 4.1e-42 | 7.4e-26 | 6.9e+03 |
| NH3 | 4.7e-17 | 1.4e+00 | 6.5e-14 | 3.6e-07 | 9.9e-06 | 7.3e-22 | 3.9e+05 | 1.9e-20 | 1.7e-15 | 2.1e-14 |
| H2S | 1.6e-02 | 9.8e+05 | 3.1e+00 | 2.8e+01 | 2.7e+02 | 2.5e-10 | 8.9e+05 | 6.1e-09 | 1.9e+03 | 1.8e+04 |
| SO2 | 5.6e-22 | 2.4e-12 | 2.2e-12 | 2.4e-12 | 2.4e-12 | 2.2e-45 | 9.8e+05 | 7.1e-40 | 1.1e-22 | 3.6e+05 |
| ClH | 5.8e-09 | 9.4e+05 | 1.5e-04 | 2.2e-04 | 1.7e+03 | 2.5e-12 | 9.0e+05 | 7.3e+02 | 6.0e+03 | 2.4e+04 |
Note. — Q1, Q2, and Q3 represent the first (25th percentile), second (median, 50th percentile), and third (75th percentile) quartiles, respectively. These data are represented in Figure 10(c,d), where interior dotted lines show Q1 and Q3 and dashed lines show Q2 (median).
Appendix C Comparison with FactSage 8.2
We independently verified the results of our model by comparison to FactSage calculations (version 8.2, Bale et al., 2016), assuming ideal gas behavior (Table C1). This comparison reveals that the partial pressure of a given volatile differs by at most a few percent. Additional comparison cases are available in the open-source code including cases with condensates.
| Case | Atmodeller | FactSage 8.2 | |||||||||||
| Hydrogen | C/H | Temp | CO | CO2 | H2 | H2O | CH4 | CO | CO2 | H2 | H2O | CH4 | |
| Oceans | kg/kg | IW† | K | bar | bar | ||||||||
| 3 | 1 | -2.0 | 1400 | 6.2 | 0.2 | 175 | 13.8 | 38.1 | 6.2 | 0.2 | 176 | 13.8 | 38 |
| 3 | 1 | +0.5 | 1400 | 46.6 | 30.7 | 237 | 337 | 28.7 | 46.4 | 30.9 | 237 | 337 | 28.7 |
| 1 | 0.1 | +2.0 | 1400 | 0.9 | 3.2 | 27.8 | 218 | 0 | 0.9 | 3.3 | 27.4 | 218 | 0 |
| 1 | 5 | +4.0 | 1400 | 9.6 | 358 | 5.4 | 432 | 0 | 10.21 | 357 | 5.8 | 432 | 0 |
| 1 | 1 | 0 | 873 | 0 | 0 | 59 | 18.3 | 19.5 | 0 | 0 | 59 | 18.3 | 19.5 |
Appendix D Bounded real gas equation of state
Real gas EOS are usually calibrated with experimental data, and are therefore valid within a certain range of temperature and pressure. Even if the parameters of a real gas EOS are within the calibrated range, numerical solvers perform better when the function is both bounded and smooth because these properties help avoid instability and ensure reliable convergence. Boundedness ensures that the solution does not experience unphysical behaviors, such as division by zero or infinite values, which can lead to numerical errors or divergence. Smoothness, on the other hand, helps prevent large, abrupt changes in the EOS that can cause solvers to struggle with step-size control or lead to inaccurate approximations, ultimately improving the robustness and accuracy of the numerical methods.
For a given EOS, a pressure range is defined based on calibration data to bracket where the EOS is physically reasonable and well behaved. Below the minimum calibration pressure, often around 1 bar, the gas species is set to obey the ideal gas law. Above the maximum calibration pressure , we let the compressibility factor (e.g., Equation 1) follow an empirical virial-like equation (e.g., Saxena & Fei, 1987a, b):
where is the gas constant, temperature, molar volume, the compressibility factor of the EOS evaluated at (, T), and the gradient of with respect to above , which could be ignored (set to zero) for simplicity. The volume integral between and can be expressed analytically
Equation D2 can be included as part of a volume integral calculation when , from which the fugacity can be determined using
Appendix E Treatment of oxygen
The condensed reservoir (molten/solid rock) of a rocky planet dominates its gaseous reservoir in terms of mass or number of moles. During the magma ocean stage, in which the molten mantle and atmosphere communicate due to similar dynamic timescales, it is therefore expected that mantle chemistry plays a role in regulating the chemistry of the combined atmosphere-mantle system. In particular, oxygen fugacity (O2) is a crucial parameter in mantle geochemical systems because it governs the oxidation state of elements, influencing mineral stability, phase equilibria, and the speciation of volatiles such as hydrogen, carbon, and sulfur. It is reasonable to assume that oxygen fugacity (O2) is buffered by the IW equilibrium in the early history of a rocky planet because planetary accretion and differentiation involve metal-silicate equilibration under reducing conditions.
Rock analyses confirm the redox state of Earth’s magma ocean was close to the IW buffer (Sossi et al., 2020). Hence, for the high-temperature calculations of both TRAPPIST-1e and K2-18b the control of mantle chemistry on atmospheric speciation is encapsulated in the constraint on O2, which is defined relative to the IW buffer that exclusively determines (absolute) O2 (in bar) at a given temperature and a reference pressure of 1 bar. Representing O2 relative to the IW buffer does not necessitate that the atmosphere of K2-18b is actually buffered by the Fe–FeO reaction at present day. Furthermore, it should be appreciated that imposing a buffered oxygen constraint does not in itself constrain the total oxygen abundance. This is evident, for example, by considering the following redox reaction:
where the equilibrium constant of the reaction is defined as
where is temperature and fugacity, in which the later can be considered equivalent to partial pressure for this discussion. If O2 is constrained relative to the IW buffer, then both O2 and are solely functions of temperature and these collectively govern the ratio of H2 to H2O but importantly not their absolute values. Imposing an additional hydrogen abundance constraint (or alternatively pressure constraint) is required to break the degeneracy to allow a unique determination of H2 and H2O. That is, two equations (Equation E3 and hydrogen mass balance) can be solved to determine two unknowns (H2 and H2O), implicitly constraining the total oxygen abundance. Hence, if O2 is prescribed as a constraint, Atmodeller solves for the total oxygen abundance self-consistently as part of the solution process. For the cooled atmosphere calculations of TRAPPIST-1e, the elemental abundance of species in the atmosphere (including oxygen) are prescribed as constraints, in which case Atmodeller then ensures a self-consistent determination of O2.
Appendix F Interplay of Solubility and Nonideality
Chachan & Stevenson (2018) and Kite et al. (2019) present models for H2 dissolution in sub-Neptunes, for ideal and nonideal scenarios, respectively. Here, we develop a toy model to demonstrate how these previous results are also obtained with the formulation of Atmodeller. We start with a simplified mass balance equation used by Atmodeller (e.g., Appendix C, Bower et al., 2022), where the first term is the dissolved mass of volatile, the second term the mass of volatile in the atmosphere, and the third term the total mass of volatile
where:
is the volatile abundance in the melt;
is the mass of melt;
is the surface radius;
is gravity, positive by definition;
is the molar mass of the volatile;
is the mean molar mass of the atmosphere;
is the surface partial pressure of the volatile; and
is the total volatile mass.
We employ a general expression for solubility with a power-law dependence on fugacity
where , , and are constants specific to a particular equilibrium between a gas species and its dissolved component and is temperature at the atmosphere-interior interface. We define a functional form for the fugacity coefficient (Equation 1) that captures the general tendency of fugacity coefficients to increase exponentially with pressure
where is total pressure, is a constant that regulates nonideality and bar is the pressure at which the volatile behaves ideally. Hence ideal behavior is recovered as or . For simplicity, temperature dependence can also be accounted for through the choice of P0. Substituting Equations F2 and F3 into F1 gives
where
and is the mass of the rocky planet, i.e. silicate mantle plus iron core. Equation F4 can be simplified for a single species atmosphere () and for , appropriate for H2 (Hirschmann et al., 2012):
The ratio of the terms in the square brackets corresponds to the ratio of the dissolved volatile mass and the atmospheric volatile mass
where is the gravitational constant. For the limit of ideal behavior () and a fully molten core () the scaling in Equation 12 of Chachan & Stevenson (2018) is recovered. Hence, the dissolved volatile mass fraction relative to that of the atmosphere increases with planet size and higher surface temperature (Figure 1, Chachan & Stevenson, 2018; Gillmann et al., 2024). Kite et al. (2019) additionally include the nonideal dissolution of H2 into magma driven by the nonlinear increase of fugacity with pressure (quantified by ), which again gives rise to enhanced volatile storage in the interior.
Appendix G Additional K2-18b Figures
Figures G1 and G2 show how atmospheric pressures (a), atmospheric C/H and C/O ratios by mass (b), solubilities (c), and fugacity coefficients (d) vary with oxygen fugacity and total C/H mass ratio, respectively, between the real and ideal cases. In these simulations, the planet mass, surface radius, and temperature are the same as those in Section 3.2 and Figures 6 and 7, and only one input parameter is varied (either oxygen fugacity or C/H mass ratio), while the others are kept constant.